OpenStudy (anonymous):

derivative using quotient rule of y=4x^3+1/2x^2-4

OpenStudy (anonymous):

\[\frac{ d }{ dx }\left( \frac{ u }{ v } \right)=\frac{ v u'-uv' }{ v ^{2} }\]